# Collapse or Thrive? Perils and Promises of Synthetic Data in a Self-Generating World

### v1

@@ Abstract
The increasing presence of AI-generated content on the internet raises a critical question: What happens when generative machine learning models are pretrained on web-scale datasets containing data created by earlier models? Some authors prophesy model collapse under a ‘replace’ scenario: a sequence of models, the first trained with real data and each later one trained only on synthetic data from its preceding model. In this scenario, models successively degrade. Others see collapse as easily avoidable; in an ‘accumulate’ scenario, a sequence of models is trained, but each training uses all real and synthetic data generated so far. In this work, we deepen and extend the study of these contrasting scenarios. First, collapse versus avoidance of collapse is studied by comparing the replace and accumulate scenarios on each of three prominent generative modeling settings; we find the same contrast emerges in all three settings. Second, we study a compromise scenario; the available data remains the same as in the accumulate scenario – but unlike accumulate and like replace, each model is trained using a fixed compute budget; we demonstrate that model test loss on real data is larger than in the accumulate scenario, but apparently plateaus, unlike the divergence seen with replace . Third, we study the relative importance of cardinality and proportion of real data for avoiding model collapse. Surprisingly, we find a non-trivial interaction between real and synthetic data, where the value of synthetic data for reducing test loss depends on the absolute quantity of real data. Our insights are particularly important when forecasting whether future frontier generative models will collapse or thrive, and our results open avenues for empirically and mathematically studying the context-dependent value of synthetic data.
1 Introduction: Model Collapse & Why It Matters
@@
With each passing day, the internet contains increasingly more AI-generated content (Altman, 2024). What is the impact of this for future of deep generative models pretrained on web-scale datasets containing data generated by their predecessors? Previous work forewarned that such model-data feedback loops can exhibit model collapse, a phenomenon whereby model performance degrades with each model-fitting iteration such that newer models trend towards useless (Shumailov et al., 2023). This prophecy is deeply concerning because society is increasingly relying on these deep generative models (Bommasani et al., 2022; Reuel et al., 2024; Perrault & Clark, 2024; Kapoor et al., 2024), and model collapse threatens that future models will be made useless as society’s current data practices pollute the pretraining data supply.
However, the model collapse literature is replete with different experimental methodologies and different mathematical assumptions of different generative models, with different papers reaching different conclusions (Taori & Hashimoto, 2023; Hataya et al., 2023; Martínez et al., 2023a; Shumailov et al., 2023; Alemohammad et al., 2023; Martínez et al., 2023b; Bohacek & Farid, 2023; Guo et al., 2023; Bertrand et al., 2023; Briesch et al., 2023; Dohmatob et al., 2024a; b; Gerstgrasser et al., 2024; Seddik et al., 2024; Marchi et al., 2024; Padmakumar & He, 2024; Chen et al., 2024; Ferbach et al., 2024; Veprikov et al., 2024). These mixed methods and findings make assessing the probability and harm of model collapse difficult.
In this work, we extend the accumulate workflow studied in Gerstgrasser et al. (2024) to cover several settings previously not studied in the literature to exhibit its effectiveness over the replace workflow. We begin by testing the following hypothesis – that model collapse emerges in a scenario where models are trained on evolving datasets built by deleting past data en masse; and model collapse is avoided in a scenario where the training datasets instead accumulate all real and synthetic data. We consider whether these claims hold in three new generative model task settings pointed to by recent prominent work; we find the claims hold. We then compare three clear dataset evolution scenarios, focusing on a new middle ground where data accumulate over time but each model is trained under a fixed compute budget; in this middle ground, we find that losses on real test data climb faster than without a compute budget but plateau to lower values than if data are deleted en masse after each model-fitting iteration. These results are consistent across five different generative modeling settings. Lastly, we investigate, in a specific situation, whether the proportion or cardinality of initial real data matters more for preventing model collapse and discover a non-trivial interaction between real and synthetic data: when real data are scarce, an appropriate amount of synthetic data reduces the test loss on real data, whereas when real data are ample, synthetic data increases the test loss on real data. Altogether, our work provides valuable comprehensive insights for predicting likely futures of deep generative models pretrained on web-scale data.
@@ 2 Testing Two Model Collapse Claims in Three New Generative Modeling Settings
Gerstgrasser et al. (2024) recently made two claims about model collapse:
@@
    Many previous papers induced model collapse by deleting past data en masse and training largely (or solely) on synthetic data from the latest generative model, and
@@
    If new synthetic data are instead added to real data, i.e., data accumulate over time, then model collapse is avoided.
These two claims are important for forecasting the future of generative models because, if correct, model collapse is then less likely to pose a realistic threat since accumulating data over time is a more realistic modeling assumption; as a partner at Andreessen Horowitz elegantly explained, deleting data en masse is “not what is happening on the internet. We won’t replace the Mona Lisa or Lord of the Rings with AI generated data, but the classics will continue to be part of the training data set" (Appenzeller, 2024).
However, these claims have not been tested in three new generative modeling settings recently introduced by prominent work (Shumailov et al., 2024) for studying model collapse:
  1. 1.
    Multivariate Gaussian Modeling: Multivariate Gaussians are repeatedly fit to data and then used to sample new synthetic data for future Gaussian fitting.
  2. 2.
    Kernel Density Estimation: Kernel density estimators are repeatedly fit to data and then used to sample new synthetic data for future kernel density estimators.
  3. 3.
    Supervised Finetuning of Language Models: Language models are finetuned in a supervised manner and then used to sample new synthetic text for future finetuning.
In this section, we ask and answer:
In these three new generative modeling settings, is model collapse caused by deleting data en masse and avoided by instead accumulating data?
In all three settings, we empirically find and (when possible) mathematically prove the answer is yes.
@@ 2.1 Model Collapse in Multivariate Gaussian Modeling
Figure 1: Model Collapse in Multivariate Gaussian Modeling. Top: Previous work (Shumailov et al., 2024) proves model collapse occurs if one iteratively fits means and covariances to data and then samples new data from a Gaussian with the fitted parameters (left). We demonstrate that if one doesn’t delete all data after each model-fitting iteration - i.e., if data accumulate - then model collapse does not occur (right). Number of Samples Per Iteration: $316$. Note: We visualize the fit Gaussians as zero-mean for easy comparison of the fit covariances across model-fitting iterations. Middle: If data are replaced, then the empirically fit means drift away from the original data’s mean with increasing model-fitting iterations, but if data instead accumulate, then the empirically fit means stabilize. Bottom: If data are replaced, then the empirically fit covariances collapse compared to the original data’s covariance, but if past data are not discarded, then the fit covariances stabilize quickly and collapse is avoided. Note: Rows 2 and 3 correspond to $d=31$ dimensional data.
We consider repeatedly fitting multivariate Gaussians to data and sampling from the fit Gaussians. We begin with $n$ real data drawn from a multivariate Gaussian with mean $\mu^{(0)}$ and covariance $\Sigma^{(0)}$:
$X_{1}^{(0)},...,X_{n}^{(0)}\sim_{i.i.d.}\mathcal{N}(\mu^{(0)},\Sigma^{(0)}).$
For model fitting, we compute the unbiased mean and covariance of the most recent data:
$\hat{\mu}_{\text{Replace}}^{(t+1)}$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{n}\sum_{j=1}^{n}X_{j}^{(t)}$ (1)
$\hat{\Sigma}_{\text{Replace}}^{(t+1)}$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{n-1}\sum_{j=1}^{n}(X_{j}^{(t)}-\hat{\mu}_{\text{Replace}}^{(t+1)})(X_{j}^{(t)}-\hat{\mu}_{\text{Replace}}^{(t+1)})^{T}$ (2)
For model sampling, we sample $m$ new synthetic data using the fit Gaussian parameters:
$X_{1}^{(t)},...,X_{n}^{(t)}\;\Big{|}\;\hat{\mu}_{\text{Replace}}^{(t)},\hat{\Sigma}_{\text{Replace}}^{(t)}\quad\sim_{i.i.d.}\quad\mathcal{N}(\hat{\mu}_{\text{Replace}}^{(t)},\hat{\Sigma}_{\text{Replace}}^{(t)}).$ (3)
Under the above data-model feedback loop, Shumailov et al. (2024) prove that
$\hat{\Sigma}_{\text{Replace}}^{(t+1)}\overset{a.s.}{\rightarrow}0\quad;\quad\mathbb{E}[\mathbb{W}_{2}^{2}(\mathcal{N}(\hat{\mu}_{\text{Replace}}^{(t+1)},\hat{\Sigma}_{\text{Replace}}^{(t+1)}),\mathcal{N}(\mu^{(0)},\Sigma^{(0)}))]\rightarrow\infty\text{ as }t\rightarrow\infty,$ (4)
where $\mathbb{W}_{2}$ denotes the Wasserstein-2 distance. This result states that the fit covariance will collapse to 00 and that the Wasserstein-2 distance will diverge as this model-data feedback loop unfolds. Note that the Wasserstein-2 distance diverges not because the covariance collapses to 00 but because the distance between the $t$-th fit mean $\hat{\mu}_{\text{Replace}}^{(t)}$ and the true mean $\mu^{(0)}$ diverges.
@@
However, this result assumes that all data are deleted after each model-fitting iteration. As discussed above, this assumption is likely unrealistic because society does not delete earlier content from the internet and replace it with new model-generated content after fitting each state-of-the-art model. What happens if data instead accumulate across model-fitting iterations? To study this, we instead consider fitting to all previous real and synthetic data:
$\hat{\mu}_{\text{Accumulate}}^{(t+1)}$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{n(t+1)}\sum_{i=0}^{t}\sum_{j=1}^{n}X_{j}^{(i)}$ (5)
$\hat{\Sigma}_{\text{Accumulate}}^{(t+1)}$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{n(t+1)-1}\sum_{i=0}^{t}\sum_{j=1}^{n}(X_{j}^{(i)}-\hat{\mu}_{\text{Accumulate}}^{(t+1)})(X_{j}^{(i)}-\hat{\mu}_{\text{Accumulate}}^{(t+1)})^{T}$ (6)
@@
Data are then sampled using these fit Accumulate parameters rather than the fit Replace parameters.
@@
Empirically, we find that deleting all data after each model-fitting iteration causes model collapse (Fig. 1 Left), whereas accumulating data across model-fitting iterations prevents model collapse (Fig. 1 Right). More specifically, we find that if data are deleted, the squared error between the fit mean $\hat{\mu}_{\text{Replace}}^{(n)}$ and the initial mean $\mu^{(0)}$ diverges (Fig. 1, Middle Left), and the fit covariance $\hat{\Sigma}_{\text{Replace}}^{(n)}$ relative to the initial covariance $\Sigma^{(0)}$ collapses to 00 (Fig. 1, Bottom Left), as measured by the ratio between the trace of $\hat{\Sigma}^{(t)}$ and the trace of $\Sigma^{(0)}$. In contrast, if data accumulate, the squared error between the fit mean and the initial mean plateaus quickly (Fig. 1, Middle Right), as does the fit covariance relative to the initial covariance (Fig. 1, Bottom Right).
Additionally, in the univariate case, we mathematically characterize the limit distribution:
@@
For notational efficiency, for a univariate Gaussian, let $\hat{\mu}^{(t)}$ and $\hat{\sigma}^{(t)}$ denote $\hat{\mu}^{(t)}_{\textrm{Accumulate}}$ and $\hat{\Sigma}^{(t)}_{\textrm{Accumulate}}$. Suppose that the mean and covariance are updated as in Eqns.  5 and 6. Then
$\mathbb{E}\left(\sigma_{t}^{2}\right)=\sigma_{0}^{2}\cdot\prod_{k=1}^{t}\left(1-\frac{1}{k^{2}n}\right)\quad$ $\xrightarrow{t\rightarrow\infty}\quad\sigma_{0}^{2}\cdot\left(\frac{\sin(\pi/\sqrt{n})}{\pi/\sqrt{n}}\right)$ (7)
$\mathbb{E}[(\mu_{t}-\mu_{0})^{2}]=\sigma_{0}^{2}\cdot\left(1-\prod_{k=1}^{t}\left(1-\frac{1}{k^{2}n}\right)\right)\quad$ $\xrightarrow{t\rightarrow\infty}\quad\sigma_{0}^{2}\cdot\left(1-\frac{\sin(\pi/\sqrt{n})}{\pi/\sqrt{n}}\right).$ (8)
See Appendix Sec. A for the proof. This reveals two key differences when data accumulate: the covariance no longer collapses, and the mean no longer diverges, meaning model collapse is mitigated.
Figure 2: Model Collapse in Kernel Density Estimation. Left: We consider 4 standard datasets from sklearn: Blobs, Circles, Moons and Swiss Roll. Center: For all four datasets, deleting data en masse causes the negative log likelihoods (NLL) of real test data to increase with each model-fitting iteration. Right: For all four datasets, accumulating data avoids model collapse. Interestingly, for specific pairs of datasets and number of samples per iteration, training on real and accumulating synthetic data can yield lower loss on real test data than training on real data alone.

### v4

@@ Abstract
What happens when generative machine learning models are pretrained on web-scale datasets containing data generated by earlier models? Some prior work warns of “model collapse” as the web is overwhelmed by synthetic data; other work suggests the problem can be contained (i.e. collapse can be avoided) by managing how available data are used in pretraining. In this paper, we report experiments on three ways of using data (training-workflows), across three generative model task-settings (multivariate Gaussian estimation, kernel density estimation, and language-model fine-tuning) to further confirm the possibility of containment: (a) we confirm that the training-workflow of replacing all real data by successive generations of purely synthetic data indeed suffers model collapse in all task-settings studied; (b) we consider the training-workflow of accumulating synthetic data alongside real data and training on all data combined and confirming that, although the proportion of real data eventually becomes zero, models remain stable and their test losses do not diverge under this training-workflow; (c) we consider a training-workflow where real and synthetic data accumulate together but successive generations of pretraining are constrained to use fixed-size data subsets each generation. In this workflow, we observe slow and gradual rather than explosive degradation of test loss performance across generations. Our insights are particularly important when forecasting whether future frontier generative models will collapse or thrive, and our results open avenues for empirically and mathematically studying the context-dependent value of synthetic data.
marginparsep has been altered.
topmargin has been altered.
marginparpush has been altered.
The page layout violates the ICML style. Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you. We’re not able to reliably undo arbitrary changes to the style. Please remove the offending package(s), or layout-changing commands and try again.
Collapse or Thrive?
Perils and Promises of Synthetic Data in a Self-Generating World
Joshua Kazdan* 1  Rylan Schaeffer* 2  Apratim Dey1  Matthias Gerstgrasser2  Rafael Rafailov2  David Donoho1  Sanmi Koyejo2
footnotetext: *Equal contribution 1Stanford Statistics 2Stanford Computer Science. Correspondence to: Joshua Kazdan <jkazdan@stanford.edu>, Rylan Schaeffer <rschaef@cs.stanford.edu>, Sanmi Koyejo <sanmi@cs.stanford.edu>.
1 Introduction
@@
With each passing day, the internet contains more AI-generated content (Altman, 2024). What impact will this have on future deep generative models and their successors? Previous work forewarned that sequences of models trained exclusively on successive iterations of their own outputs can exhibit model collapse: model performance degrades with each iteration and later models become progressively more unusable (Shumailov et al., 2023). Due to increasingly heavy usage of LMs to produce written content for the internet (Bommasani et al., 2022; Reuel et al., 2024; Perrault & Clark, 2024; Kapoor et al., 2024), model collapse alarmists believe current trends, if unmitigated, will irredeemably pollute the pretraining data supply and degrade language models across time.
However, the model collapse literature contains a variety of methodologies, assumptions, and models. Research supports conflicting positions about such future scenarios (Taori & Hashimoto, 2023; Hataya et al., 2023; Martínez et al., 2023; Shumailov et al., 2023; Alemohammad et al., 2024; Martínez et al., 2023; Bohacek & Farid, 2023; Guo et al., 2024; Bertrand et al., 2024; Briesch et al., 2023; Dohmatob et al., 2024a; b; Gerstgrasser et al., 2024; Seddik et al., 2024; Marchi et al., 2024; Padmakumar & He, 2024; Chen et al., 2024; Ferbach et al., 2024a; Veprikov et al., 2024; Dey & Donoho, 2024; Gillman et al., 2024). Concordance in the field is especially challenging because “model collapse” has been defined in various ways and studied under different assumptions concerning training-workflows and task-settings. Shumailov et al. (2023) defined model collapse as a “degenerative process affecting generations of learned generative models, in which the data they generate end up polluting the training set of the next generation.” Dohmatob et al. (2024b) called model collapse the worsening of scaling law curves when training on synthetic as opposed to real data. In their theory sections, Shumailov et al. (2024) and Gerstgrasser et al. (2024) defined model collapse as divergent test loss after multiple iterations of training in certain task settings. Due to space constraints, we discuss Related Work more extensively in Appendix A.
In this paper, we take model collapse by its literal meaning: that model performance deteriorates to the level of uselessness across iterations of an assumed data training-workflow under an assumed generative model task-setting. We bring new clarity to the discussion by studying how three different data training-workflows shape trained model performance over multiple generations. The first training-workflow, initially studied by Shumailov et al. (2023), fully replaces previous data at each iteration with newly generated synthetic data from the most recent generative model. In particular, under the replace training-workflow, real data are only used at the very first iteration. We conduct experiments in three different generative model task-settings - from basic multivariate Gaussian modeling (MGM) to kernel density estimation (KDE) to supervised fine-tuning of language models (SFT) – and observe that in each setting, the replace training-workflow induces collapse. We also consider the data training-workflow of accumulating all prior data - both real and synthetic - and find that on all three task-settings, collapse does not occur: the test loss on real data stays bounded at reasonable levels. Additionally, we introduce and study a more realistic “fixed compute” training-workflow. In this training-workflow, all real and synthetic data are retained, as in the accumulate setting. However, a dataset is subsampled from the total available pool of data to train the next model. While the pool of data grows over time, the size of the subsampled training dataset remains constant regardless of the model iteration. This training-workflow reflects the reality that available compute might not keep pace with the rate of synthetic data generation, forcing model providers to sample a fraction of future web-scale data for training. This training-workflow can be viewed as a middle ground between the assumed fixed-sample size replace training-workflow and the increasing sample-size accumulate training-workflow. We find that test errors grow larger and more quickly than in the accumulate scenario, but more slowly than they grow in the replace scenario. These results are consistent across five different generative model task-settings. Third, we investigate whether the proportion or cardinality of initial real data matters more for preventing model collapse and discover a non-trivial interaction between real and synthetic data: when real data are scarce, adding a small amount of synthetic data can reduce test loss, whereas when real data are ample, any synthetic data increases the test loss calculated on real data.
To summarize our contributions:
  1. 1.
    We show that the hypotheses of Gerstgrasser et al. (2024) are more universally applicable than originally claimed: three generative model settings widely cited as evidence for the dangers of collapse under a replace training-workflow (Shumailov et al., 2024) fail to collapse under the accumulate training-workflow.
  2. 2.
    We characterize test-loss behavior in a middle-ground between the replace and accumulate workflows. This middle ground reflects the impact of limited training-compute in a data ecosystem that is being flooded with synthetic data.
  3. 3.
    We illustrate that when real data are in limited supply, supplementing the training set by adding small amounts of synthetic data can improve test loss. On the other hand, when real data are abundant, supplementation with synthetic data only increases test loss.
Figure 1: Model Collapse in Multivariate Gaussian Modeling. Top: Previous work (Shumailov et al., 2023; Alemohammad et al., 2024; Bertrand et al., 2024) proved model collapse occurs under the replace training-workflow which iteratively fits means and covariances to data, deletes earlier data, and replaces it with samples from a Gaussian with the fitted parameters (left). However, under the accumulate workflow where one doesn’t delete data after each model-fitting iteration, model collapse does not occur (right). Note: We visualize the fit Gaussians as zero-mean for easy comparison of the fit covariances across model-fitting iterations. Middle: If data are replaced, then the fitted means drift away from the original data’s mean, but if data instead accumulate, then the fitted means stabilize. Bottom: If data are replaced, then the fitted covariances collapse compared to the original data’s covariance, but if past data are not discarded, the fitted covariances stabilize quickly and collapse is averted.
@@ 2 Testing Two Model Collapse Claims in Three New Generative Modeling Settings
Gerstgrasser et al. (2024) recently made two claims:
@@
    Where model collapse has been documented, the collapse can be attributed to the replace workflow, where at each iteration, prior data (synthetic or real) are deleted en masse.
@@
    In the accumulate workflow where synthetic data accumulate alongside the original real data, model collapse is avoided.
If correct, these two claims suggest that model collapse is less likely to pose a significant threat to future deep generative models since accumulating data over time is a more realistic model of internet evolution. As a partner at Andreessen-Horowitz elegantly explained, deleting data en masse is “not what is happening on the internet. We won’t replace the Mona Lisa or Lord of the Rings with AI-generated data, but the classics will continue to be part of the training data set” (Appenzeller, 2024). Despite such plausible heuristics, the claims of Gerstgrasser et al. (2024) were not formally tested in three generative modeling task-settings studied in influential recent work (Shumailov et al., 2024): (1) Multivariate Gaussian Modeling (MGM); (2) Kernel Density Estimation (KDE); and (3) Supervised Fine-tuning of Language Models (SFT). In this section, we ask and answer: In these three new generative modeling settings, do Gerstgrasser et al. (2024) Claims 1 and 2 continue to hold? In all three settings, we find that the answer is yes.
@@ 2.1 Model Collapse in Multivariate Gaussian Modeling
Figure 2: Model Collapse in Kernel Density Estimation. Left: We consider 4 standard datasets from sklearn: Blobs, Circles, Moons and Swiss Roll. Center: For all four datasets, deleting data en masse causes the negative log likelihoods (NLL) of held-out real data to increase with each model-fitting iteration. Right: For all four datasets, the accumulate training-workflow avoids diverging test loss on real data. Interestingly, for specific pairs of datasets and numbers of samples per iteration, training on real and accumulated synthetic data can yield lower test-loss on held-out real data than would training on real data alone.
Following Shumailov et al. (2023); Alemohammad et al. (2024); Bertrand et al. (2024), we study what happens when one iteratively fits multivariate Gaussians and samples from them. We begin with $n$ real data drawn from a Gaussian with mean $\mu^{(0)}$ and covariance $\Sigma^{(0)}$: $X_{1}^{(0)},...,X_{n}^{(0)}\sim_{i.i.d.}\mathcal{N}(\mu^{(0)},\Sigma^{(0)}).$ For model fitting, we compute the unbiased mean and covariance of the most recent data:
$\hat{\mu}_{\text{Replace}}^{(t+1)}$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{n}\sum_{j=1}^{n}X_{j}^{(t)}$
$\hat{\Sigma}_{\text{Replace}}^{(t+1)}$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{n-1}\sum_{j=1}^{n}(X_{j}^{(t)}-\hat{\mu}_{\text{Replace}}^{(t+1)})(X_{j}^{(t)}-\hat{\mu}_{\text{Replace}}^{(t+1)})^{T}$
For sampling, we draw $n$ synthetic data using the fit parameters: $X_{1}^{(t)},...,X_{n}^{(t)}\quad\sim_{i.i.d.}\quad\mathcal{N}(\hat{\mu}_{\text{Replace}}^{(t)},\hat{\Sigma}_{\text{Replace}}^{(t)}).$ Shumailov et al. (2024) proved that as $t\rightarrow\infty$,
$\hat{\Sigma}_{\text{Replace}}^{(t+1)}$ $\overset{a.s.}{\rightarrow}0$
[…]
@@
where $\mathbb{W}_{2}$ denotes the Wasserstein-2 distance. This result simply states that the fit covariances will collapse to 00 and that the Wasserstein-2 distance will diverge as this model-data feedback loop unfolds. However, this result assumes that all data are replaced with new synthetic data after each model-fitting iteration, which is likely unrealistic as a model of data evolution; as a rule we do not delete earlier content from the internet, replacing it en masse with new model-generated content after fitting each state-of-the-art model. To study what happens if data instead accumulate across model-fitting iterations, we fit to all previous real and synthetic data, a mixture in which the fraction of real data asymptotically approaches 00:
$\hat{\mu}_{\text{Accumulate}}^{(t+1)}$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{n(t+1)}\sum_{i=0}^{t}\sum_{j=1}^{n}X_{j}^{(i)}$
$\hat{\Sigma}_{\text{Accumulate}}^{(t+1)}$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{n(t+1)-1}\sum_{i=0}^{t}\sum_{j=1}^{n}\bar{X}_{j}^{(i)}(\bar{X}_{j}^{(i)})^{T},$
@@
where $\bar{X}_{j}^{(i)}\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}X_{j}^{(i)}-\hat{\mu}_{\text{Accumulate}}^{(t+1)}$ is shorthand for the centered datum. Data are then sampled using these fit Accumulate parameters rather than the fit Replace parameters.
Figure 3: Model Collapse in Supervised Fine-tuning of Language Models. Fine-tuning Google’s Gemma2 models on Nvidia’s HelpSteer 2 dataset demonstrates that model collapse occurs if previous data are replaced after each model-fitting iteration (left), whereas model collapse is avoided if new synthetic data instead accumulate with previous real and synthetic data (right).
@@
Empirically, we find that replacing all data with new synthetic data after each model-fitting iteration causes model collapse (Fig. 1 Left), whereas accumulating data across model-fitting iterations prevents model collapse (Fig. 1 Right). More specifically, we find that if data are deleted, the squared error between the fit mean $\hat{\mu}_{\text{Replace}}^{(n)}$ and the initial mean $\mu^{(0)}$ diverges (Fig. 1, Middle Left), and the fit covariance $\hat{\Sigma}_{\text{Replace}}^{(n)}$ relative to the initial covariance $\Sigma^{(0)}$ collapses to 00 (Fig. 1, Bottom Left). In contrast, if data accumulate, the squared error between the fit mean and the initial mean plateaus quickly (Fig. 1, Middle Right), as does the fit covariance relative to the initial covariance (Fig. 1, Bottom Right). For univariate Gaussians, we can mathematically characterize the limit distribution:
@@
For notational efficiency, for a univariate Gaussian, let $\hat{\mu}^{(t)}$ and $\hat{\sigma}^{(t)}$ denote $\hat{\mu}^{(t)}_{\textrm{Accumulate}}$ and $\hat{\Sigma}^{(t)}_{\textrm{Accumulate}}$. Then
$\mathbb{E}\left[\sigma_{t}^{2}\right]\quad$ $\xrightarrow{t\rightarrow\infty}\quad\sigma_{0}^{2}\cdot\left(\frac{\sin(\pi/\sqrt{n})}{\pi/\sqrt{n}}\right)$ (1)
$\mathbb{E}[(\mu_{t}-\mu_{0})^{2}]\quad$ $\xrightarrow{t\rightarrow\infty}\quad\sigma_{0}^{2}\cdot\left(1-\frac{\sin(\pi/\sqrt{n})}{\pi/\sqrt{n}}\right).$ (2)
See Appendix B for the proof. This reveals two key differences when data accumulate: the covariance no longer collapses, and the mean no longer diverges. This also implies that the Wasserstein-2 distance no longer diverges. Altogether, if data accumulate, model collapse is mitigated.